# Adversarial Vulnerabilities in Large Language Models for Time Series Forecasting

### v3

@@ Abstract
Large Language Models (LLMs) have recently demonstrated significant potential in time series forecasting, offering impressive capabilities in handling complex temporal data. However, their robustness and reliability in real-world applications remain under-explored, particularly concerning their susceptibility to adversarial attacks. In this paper, we introduce a targeted adversarial attack framework for LLM-based time series forecasting. By employing both gradient-free and black-box optimization methods, we generate minimal yet highly effective perturbations that significantly degrade the forecasting accuracy across multiple datasets and LLM architectures. Our experiments, which include models like LLMTime with GPT-3.5, GPT-4, LLaMa, and Mistral, TimeGPT, and TimeLLM show that adversarial attacks lead to much more severe performance degradation than random noise, and demonstrate the broad effectiveness of our attacks across different LLMs. The results underscore the critical vulnerabilities of LLMs in time series forecasting, highlighting the need for robust defense mechanisms to ensure their reliable deployment in practical applications.
@@ 1 Introduction
Time series forecasting plays a pivotal role in numerous real-world applications, ranging from finance and healthcare to energy management and climate modeling. Accurately predicting temporal patterns in the data is crucial for informed decision-making in these domains [1]. Recently, Large Language Models (LLMs), originally designed for Natural Language Processing (NLP) tasks, have demonstrated remarkable potential in handling time series forecasting challenges [2, 3, 4, 5, 6]. These models, including BERT [7], GPT [8, 9], LLaMa [10] and their successors, leverage their powerful attention mechanisms and vast pre-training on diverse datasets to capture intricate temporal dependencies, making them highly effective for complex forecasting tasks.
LLMs exhibit strong generalization capabilities across various types of time series data. Compared to traditional models like ARIMA[11] and Exponential Smoothing [12], as well as advanced deep learning models such as DNNs [13, 14, 15], and Transformer-based architectures[16, 17, 18, 19], LLMs excel in modeling long-term dependencies and capturing non-linear patterns within temporal sequences. This has resulted in impressive forecasting accuracy across applications ranging from energy consumption predictions to weather forecasting [20, 21].
However, despite their success, the robustness and reliability of LLMs in real-world forecasting remain concerns, particularly their vulnerability to adversarial attacks is under-explored. Adversarial attacks introduce subtle, often imperceptible perturbations to input data, leading to significant and misleading changes in model predictions. While the susceptibility of machine learning models to such attacks has been well-explored in image processing and NLP domains [22, 23, 24], there is a noticeable gap in research on their impact on LLMs used for time series forecasting.
@@
While adversarial attacks and defenses for deep neural networks have been extensively studied across various domains [25], executing adversarial attacks against LLMs in time series forecasting presents two significant challenges. First, to prevent information leakage, we cannot use ground truth values (i.e., future time steps) when attacking forecasting models. Second, LLMs must be treated as strict black-box systems due to the difficulty of accessing their internal workings and parameters.
@@
In conclusion, this study contributes to the growing discourse on LLMs robustness by exposing their vulnerabilities to adversarial attacks in Time Series Forecasting. Our results highlight the necessity of addressing these vulnerabilities to advance the development of LLMs that are not only accurate but also resilient, thereby enhancing their practical utility in high-stakes environments.
@@ 2 Related Work
2.1 Adversarial Attacks in Time Series Forecasting
Adversarial attacks in time series forecasting have emerged as a crucial area of research, exposing vulnerabilities in forecasting models. Unlike adversarial studies in static domains, such as object recognition or time series classification, adversarial attacks on time series forecasting cannot leverage ground truth data for perturbation generation due to the risk of information leakage [26]. To address this challenge, surrogate techniques have been adopted [27], which bypass the need for labels, as is done in traditional adversarial attack methods like the Fast Gradient Sign Method [28]. Several studies have treated forecasting models as white-box systems to investigate the effects of adversarial attacks on commonly used models in time series forecasting, such as ARIMA, LSTMs, and Transformer-based models [29, 30]. These studies demonstrate that even small perturbations can severely impact these models, resulting in inaccurate forecasts. However, evaluating the vulnerability of LLM-based forecasting presents a significant challenge, as internal access is typically restricted, requiring these models to be treated as black-box systems.
@@ 2.2 Adversarial Attacks on LLMs
  • Jailbreak Prompting [31, 32]: Crafted prompts that bypass LLM guardrails, inducing unintended or harmful outputs by exploiting unconventional phrasing.
  • Prompt Injection [33, 34, 35]: Adversarial instructions embedded into benign prompts to manipulate LLM responses, highlighting their vulnerability to prompt manipulation.
  • Gradient Attacks [25, 36]: Using internal model parameters, attackers apply gradient-based methods to perturb inputs, significantly altering outputs with minimal changes.
  • Embedding Perturbations [37, 38]: Subtle changes to input embeddings disrupt LLM’s internal representations, leading to erroneous outputs with minimal visible input alterations.
@@ 3 Manipulating LLM-based Time Series Forecasting
3.1 Formulations of LLM-based Time Series Forecasting
LLMs have shown promising performance in time series forecasting by leveraging their ability to perform next-token prediction, a technique originally developed for text-based tasks [2, 6]. A typical LLM-based time series forecasting model, denoted as $f(\cdot)$, consists of two primary components: an embedding or tokenization module that encodes the time series data into a sequence of tokens, and a pre-trained LLM that autoregressively predicts the subsequent tokens. The embedding module translates the raw time series into a format suitable for the LLM, while the LLM captures the temporal dependencies and generates predictions based on its learned representations.
@@ 3.2 Threat model
In practical applications, accessing the full set of detailed parameters of an LLM is typically infeasible, which forces the attacker to treat the target model as a black-box system. Additionally, acquiring the complete training dataset is impractical, meaning the attacker lacks access to this information as well. The attacker’s capabilities can be summarized as follows: (i) no access to the training data, (ii) no access to internal information of the LLM-based forecasting model, (iii) no access to ground truth, and (iv) the ability to query the target model.
@@ 5.1 Datasets
  • ETTh1 and ETTh2 (Electricity Transformer Temperature Hourly) [16]: These datasets consist of two years of hourly recorded data from electricity transformers, capturing temperature and power consumption variables.
  • IstanbulTraffic[2]: This dataset contains hourly measurements of road traffic volumes across different sensors. It captures temporal dependencies related to traffic patterns, which are dynamic and fluctuating time series.
  • Weather [39]: This dataset comprises meteorological data, including variables such as temperature, humidity, and wind speed, recorded hourly. It provides a challenging forecasting task due to the inherent variability and complexity of weather patterns.
  • Exchange [40]: This dataset consists of daily exchange rates from eight foreign countries—Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore—covering the period from 1990 to 2016.
@@ 5.2 Target Models
  • TimeGPT [41]: A large model specifically pre-trained with a vast amount of time series data. It uses advanced attention mechanisms and temporal encoding to capture complex patterns in sequential data, making it a leading LLM designed explicitly for time series forecasting. Its pre-training, which is conducted from scratch using vast amounts of time series data, allows it to serve as a robust and versatile tool for a wide range of time-dependent applications.
  • LLMTime [2]: This model treats time series forecasting as a next-token prediction task, using LLM architectures like GPT and LLaMa. By converting time series data into numerical sequences, LLMTime enables these models to apply their sequence prediction strengths to time series. To test the robustness of our adversarial attacks, we experimented with base models including GPT-3.5, GPT-4, LLaMa, and Mistral, assessing their resilience when adapted from natural language processing to time series forecasting.
  • TimeLLM [6]: This model presents a novel approach for time series forecasting by adapting LLMs by reprogramming input time series into textual representations that are more compatible with LLMs, allowing the models to perform time series forecasting tasks without altering their pre-trained structures. The key innovation is the Prompt-as-Prefix (PaP) technique, which augments input context to guide the LLM in transforming reprogrammed data into accurate forecasts.
  • TimeNet[39] and iTransformer[18]are two supervised forecasting models that leverage the attention-based architecture of transformers, enabling them to capture long-term dependencies in time series data effectively. These models provide strong performance in time series forecasting tasks, and in this study, they are used in the performance and robustness comparison between LLM-based forecasting models with non-LLM models.
@@ 5.4 Overall Comparison
Across all models and datasets, the adversarial perturbations induced much greater disruptions than GWN, clearly impacting the predictions and demonstrating the precision of the attack in destabilizing LLM-based forecasting. This underlines the importance of developing robust defensive strategies to protect LLMs against such targeted adversarial attacks, as their current vulnerability poses a significant challenge for practical applications.
@@ 5.6 Hyperparameter Study
We analyze the impact of varying scale ratios on model performance under both GWN and DGA adversarial attacks, with the vertical axis in Figure LABEL:fig:hyper representing the increase in MAE. This experiment was conducted across three different datasets using three LLM-based forecasting models. As demonstrated in the figure, DGA consistently results in a more significant increase in MAE compared to GWN as the scale ratio rises, indicating that DGA is more effective in disrupting the model’s predictions. To balance imperceptibility and manipulation effectiveness, the perturbation scale can be chosen as 2% of the mean value of the given data.

### v4

@@ Abstract
Large Language Models (LLMs) have recently demonstrated significant potential in time series forecasting, offering impressive capabilities in handling complex temporal data. However, their robustness and reliability in real-world applications remain under-explored, particularly concerning their susceptibility to adversarial attacks. In this paper, we introduce a targeted adversarial attack framework for LLM-based time series forecasting. By employing both gradient-free and black-box optimization methods, we generate minimal yet highly effective perturbations that significantly degrade the forecasting accuracy across multiple datasets and LLM architectures. Our experiments, which include models like LLMTime with GPT-3.5, GPT-4, LLaMa, and Mistral, TimeGPT, and TimeLLM show that adversarial attacks lead to much more severe performance degradation than random noise, and demonstrate the broad effectiveness of our attacks across different LLMs. The results underscore the critical vulnerabilities of LLMs in time series forecasting, highlighting the need for robust defense mechanisms to ensure their reliable deployment in practical applications. The code repository can be found at Johnson/AdvAttackLLM4TS.
@@ 1 INTRODUCTION
Time series forecasting plays a pivotal role in numerous real-world applications, ranging from finance and healthcare to energy management and climate modeling. Accurately predicting temporal patterns in the data is crucial for informed decision-making in these domains (Liu et al., 2022b; Jiang et al., 2024). Recently, Large Language Models (LLMs), originally designed for Natural Language Processing (NLP) tasks, have demonstrated remarkable potential in handling time series forecasting challenges (Gruver et al., 2024; Liu et al., 2024b, a; Tan et al., 2024; Jin et al., 2023a). These models, including BERT (Devlin, 2018), GPT (Brown, 2020; Achiam et al., 2023), LLaMa (Touvron et al., 2023) and their successors, leverage their powerful attention mechanisms and vast pre-training on diverse datasets to capture intricate and non-linear temporal dependencies, making them highly effective for complex forecasting tasks.
LLMs exhibit strong generalization capabilities across various types of time series data. Compared to traditional models like ARIMA (Kalpakis et al., 2001) and Exponential Smoothing (Gardner Jr, 1985), as well as advanced deep learning models such as DNNs (Salinas et al., 2020; Oreshkin et al., 2019; Challu et al., 2023), and Transformer-based architectures (Zhou et al., 2021; Wu et al., 2021; Liu et al., 2024d; Zhou et al., 2022), LLMs excel in modeling long-term dependencies and capturing non-linear patterns within temporal sequences. This has resulted in impressive forecasting accuracy across applications ranging from energy consumption predictions to weather forecasting (Jin et al., 2023b).
However, despite their success, the robustness and reliability of LLMs in real-world forecasting applications remain concerns, particularly their vulnerability to adversarial attacks is under-explored. Adversarial attacks introduce subtle, often imperceptible perturbations to input data, leading to significant and misleading changes in model predictions. While the susceptibility of machine learning models to such attacks has been well-explored in image processing and NLP domains (Xu et al., 2020; Morris et al., 2020; Wei et al., 2018), there is a noticeable gap in research on their impact on LLMs used for time series forecasting.
@@
While adversarial attacks and defenses for deep neural networks have been extensively studied across various domains (Madry, 2017), executing adversarial attacks against LLMs in time series forecasting presents two significant challenges. First, to prevent information leakage, we cannot use ground truth values (i.e., future time steps) when attacking forecasting models. Second, LLMs must be treated as strict black-box systems due to the difficulty of accessing their internal workings and parameters.
@@
In conclusion, this study contributes to the ongoing discourse on the robustness of LLMs by revealing their vulnerabilities to adversarial attacks in time series forecasting. Our findings underscore the critical need to address these vulnerabilities to ensure that LLM-based forecasting models are not only accurate but also resilient, thereby enhancing their practical utility in high-stakes applications.
@@ 2 RELATED WORK
2.1 Adversarial Attacks on Forecasting
Adversarial attacks in time series forecasting have emerged as a crucial area of research, exposing vulnerabilities in forecasting models. Unlike adversarial studies in static domains, such as object recognition or time series classification, adversarial attacks on time series forecasting cannot leverage ground truth data for perturbation generation due to the risk of information leakage (Liu et al., 2022a). To address this challenge, surrogate techniques have been adopted (Liu et al., 2021), which bypass the need for labels, as is done in traditional adversarial attack methods like the Fast Gradient Sign Method (Goodfellow et al., 2014). Several studies have treated forecasting models as white-box systems to investigate the effects of adversarial attacks on commonly used models in time series forecasting, such as ARIMA, LSTMs, and Transformer-based models (Liu et al., 2022c, 2023). These studies demonstrate that even small perturbations can severely impact these models, resulting in inaccurate forecasts. However, evaluating the vulnerability of LLM-based forecasting presents a significant challenge, as internal access is typically restricted, requiring these models to be treated as black-box systems.
@@ 2.2 Adversarial Attacks on LLMs
  • Jailbreak Prompting (Yu et al., 2024; Wei et al., 2024): Crafted prompts that bypass LLM guardrails, inducing unintended or harmful outputs by exploiting unconventional phrasing.
  • Prompt Injection (Greshake et al., 2023; Xue et al., 2024; Liu et al., 2024c): Adversarial instructions embedded into benign prompts to manipulate LLM responses, highlighting their vulnerability to prompt manipulation.
  • Gradient Attacks (Madry, 2017; Guo et al., 2021): Using internal model parameters, attackers apply gradient-based methods to perturb inputs, significantly altering outputs with minimal changes.
  • Embedding Perturbations (Schwinn et al., 2024; Singh et al., 2024): Subtle changes to input embeddings disrupt LLM’s internal representations, leading to erroneous outputs with minimal visible input alterations.
@@ 3 MANIPULATING LLM-BASED TIME SERIES FORECASTING
3.1 Formulations of LLM-based Forecasting
LLMs have shown promising performance in time series forecasting by leveraging their ability to perform next-token prediction, a technique originally developed for text-based tasks (Gruver et al., 2024; Jin et al., 2023a). A typical LLM-based time series forecasting model, denoted as $f(\cdot)$, consists of two primary components: an embedding or tokenization module that encodes the time series data into a sequence of tokens, and a pre-trained LLM that autoregressively predicts the subsequent tokens. The embedding module translates the raw time series into a format suitable for the LLM, while the LLM captures the temporal dependencies and generates predictions based on its learned representations.
@@ 3.2 Threat Model
In practical applications, accessing the full set of detailed parameters of an LLM is typically infeasible, which forces the attacker to treat the target model as a black-box system. Additionally, acquiring the complete training dataset is impractical, meaning the attacker lacks access to this information as well. The attacker’s capabilities can be summarized as follows:
  • no access to the training data,
  • no access to internal information of the LLM-based forecasting model,
  • no access to ground truth,
  • the ability to query the target model.
@@ 5.1 Datasets
  • ETTh1 and ETTh2 (Electricity Transformer Temperature Hourly) (Zhou et al., 2021): These datasets consist of two years of hourly recorded data from electricity transformers, capturing temperature and power consumption variables.
  • IstanbulTraffic (Gruver et al., 2024): This dataset contains hourly measurements of road traffic volumes across different sensors. It captures temporal dependencies related to traffic patterns, which are dynamic and fluctuating time series.
  • Weather (Wu et al., 2023): This dataset comprises meteorological data, including variables such as temperature, humidity, and wind speed, recorded hourly. It provides a challenging forecasting task due to the inherent variability and complexity of weather patterns.
  • Exchange (Lai et al., 2018): This dataset consists of daily exchange rates from eight foreign countries—Australia, the United Kingdom, Canada, Switzerland, China, Japan, New Zealand, and Singapore—covering the period from 1990 to 2016.
@@ 5.2 Target Models
  • TimeGPT (Garza and Mergenthaler-Canseco, 2023): A large model specifically pre-trained with a vast amount of time series data. It uses advanced attention mechanisms and temporal encoding to capture complex patterns in sequential data, making it a leading LLM designed explicitly for time series forecasting. Its pre-training, which is conducted from scratch using vast amounts of time series data, allows it to serve as a robust and versatile tool for a wide range of time-dependent applications.
  • LLMTime (Gruver et al., 2024): This model treats time series forecasting as a next-token prediction task, using LLM architectures like GPT and LLaMa. By converting time series data into numerical sequences, LLMTime enables these models to apply their sequence prediction strengths to time series. To test the robustness of our adversarial attacks, we experimented with base models including GPT-3.5, GPT-4, LLaMa, and Mistral, assessing their resilience when adapted from natural language processing to time series forecasting.
  • TimeLLM (Jin et al., 2023a): This model presents a novel approach for time series forecasting by adapting LLMs by reprogramming input time series into textual representations that are more compatible with LLMs, allowing the models to perform time series forecasting tasks without altering their pre-trained structures. The key innovation is the Prompt-as-Prefix (PaP) technique, which augments input context to guide the LLM in transforming reprogrammed data into accurate forecasts.
  • TimeNet (Wu et al., 2023) and iTransformer (Liu et al., 2024d)are two supervised forecasting models that leverage the attention-based architecture of transformers, enabling them to capture long-term dependencies in time series data effectively. These models provide strong performance in time series forecasting tasks, and in this study, they are used in the performance and robustness comparison between LLM-based forecasting models with non-LLM models.
@@ 5.4 Overall Comparison
Across all models and datasets, the adversarial perturbations introduced significantly greater disruptions than GWN, clearly impacting the predictions and demonstrating the precision of the attack in destabilizing LLM-based forecasting. The magnitude of the degradation in predictive accuracy highlights the effectiveness of the proposed DGA. These findings emphasize the urgent need for robust defensive strategies to safeguard LLM-based forecasting models against adversarial threats. The current vulnerability of these models presents a significant challenge for real-world applications, particularly in high-stakes domains such as financial forecasting, energy demand prediction, and intelligent transportation systems. Without adequate defenses, adversarial attacks could lead to erroneous predictions, resulting in potential financial losses, inefficient resource allocations, or compromised safety in critical infrastructure.
@@ 5.6 Hyperparameter Study
We systematically analyze the impact of varying scale ratios on model performance under both GWN and DGA adversarial perturbations. The vertical axis in Figure LABEL:fig:hyper represents the increase in MAE, serving as a measure of the extent to which each attack degrades the precision of the forecast. This experiment is conducted across three distinct datasets using three LLM-based forecasting models.
As illustrated in Figure LABEL:fig:hyper, the DGA consistently induces a more pronounced increase in MAE compared to GWN as the scale ratio increases. This trend highlights the greater effectiveness of the DGA in destabilizing model predictions, as it is specifically designed to exploit model vulnerabilities rather than introduce random noise. To achieve a balance between imperceptibility and attack effectiveness, we determine that an optimal perturbation scale can be set at 2% of the mean value of the given dataset. This choice ensures that the adversarial perturbation remains subtle enough to evade detection while still significantly impairing the forecasting model’s performance.